A customer received change worth $18.95. He receives 10 more dollar coins than 25cent peices and 22 more 10cent pieces than 25cent pieces. how many coins of each did he receive?
Answer by scott8148(6628)

let x="25 cent pieces", so x+10="dollar coins" and x+22="10 cent pieces"

working in cents to avoid decimal points

x(25)+(x+10)(100)+(x+22)(10)=1895 __ 25x+100x+1000+10x+220=1895 __ 135x+1220=1895

135x=675 __ x=5